# Efficient high-resolution refinement in cryo-EM with stochastic gradient descent

### v2

@@ 4 Numerical experiments
(b) $z=0$ slice of $H$
Figure 3: The condition number $\kappa(H)$ for the experimental dataset used in Section 4, with $N=30000$ images, grid of dimensions $192\times 192\times 192$, and $\lambda=10^{-8}$. (a) Lower bound of $\kappa(H)$ (blue, solid) computed using (21), the value of $\kappa(H)$ for the given dataset without CTF (red, dashed) and with CTF (green, dashdot). The condition numbers computed for this plot are for $H$ with nearest-neighbor interpolation. (b) The $z=0$ plane of the diagonal of the Hessian matrix $H$ for the given dataset, computed using trilinear interpolation and reshaped as a 3D volume. Note the effect of the CTF, in contrast to Figure 1(c), where the particle images are not corrupted by the CTF.
@@
Figure 4: Results of the numerical experiments with trilinear interpolation projection matrices $P_{\phi}^{tri}$ for $N=30000$ images of $192\times 192$ pixels. Panel (a) shows the loss function value and panel (b) shows the FSC of the final reconstruction with a ground-truth solution for SGD with no preconditioner (blue), SGD with a precomputed preconditioner (red), and Algorithm 1, namely SGD with an estimated preconditioner (green).
@@ 5 Outlook and conclusion
    A critical component of the current EM approaches is the $\ell_{2}$ regularizer, which makes the maximization step computationally tractable. SGD, on the other hand, is compatible with other regularization methods, and one could take advantage, for example, of the Wilson prior [gilles_molecular_2022] and learned regularization methods [kimanius_exploiting_2021, kamilov_plug-and-play_2023, kimanius_data-driven_2023].
@@ 6 Acknowledgments
This work was supported by the grants NIH/R01GM136780 and AFOSR/FA9550-21-1-0317, and by the Simons Foundation. The work was done while BT was affiliated with Yale University.

### v3

@@ 4.2 Results
In Figure 7, we show a qualitative comparison between the outputs of the three SGD algorithms and the ground truth volume to illustrate the effect of the preconditioner on the high-resolution details in the reconstructions. The figure shows four different views of the volumes (one view on each row), where the volumes are displayed as the isosurface at six standard deviations above the mean voxel value of each volume. Firstly, we see that the main differences between the ground truth L-BFGS volume (first column, grey) and the volume obtained using SGD with no preconditioner (second column, blue) are in the high-resolution details, where the former shows more defined fine-scale features than the latter. Moreover, both volumes obtained using the preconditioned SGD (third column, red for the precomputed preconditioner and fourth column, green for the estimated preconditioner using Algorithm 1) show a similar level of high-resolution detail to the ground truth volume. Secondly, there is no significant visual difference between the volumes obtained using the precomputed and estimated preconditioners. These conclusions drawn from Figure 7 are consistent with the quantitative results shown in Figures 4-6, especially regarding the convergence of the SGD algorithms at low resolution (where the preconditioner does not have a significant effect) and at high resolution (where the preconditioner leads to the differences between the reconstructions), as shown in Figure 5 and Figure 6.
The aim of the figures in this section is to show, either quantitatively (Figures 4-6) or qualitatively (Figure 7), the differences between the resulting volumes and how the preconditioner leads to a different reconstruction that is closer to the numerical ground truth obtained by solving the optimization problem (12) to higher accuracy. The results in this section (and more generally in this paper) do not refer to a measure of the overall quality of a cryo-EM map, as this is dependent on multiple factors, both in terms of how the map is obtained and in terms of how such a measure of the map quality is computed. In particular, obtaining a “good” map involves multiple steps, including estimating the noise in the images, accurately estimating the pose variables, appropriately regularizing and filtering the maps (i.e. by splitting the data into half sets, computing independent reconstructions, and obtaining a Wiener filter from the FSC between the two reconstructions, that is then applied to the final map) and solving the volume reconstruction problem to high accuracy. Measuring the quality of a cryo-EM map, likewise, involves computing the FSC between independent half set reconstructions and using the appropriate FSC threshold to determine the final resolution (and hence the quality) of the map. As this paper only addresses the volume reconstruction problem (given fixed values of the other parameters such as pose variables, noise level and regularization parameter), the results in the current section only show the contribution of the preconditioner to the accuracy of the solution to the volume reconstruction problem (12) and not how the map quality would improve if such a preconditioner was used in the full cryo-EM image processing pipeline that also estimates the other parameters together with accurately measuring the final resolution based on half-set FSC curves. Thus, the usual FSC threshold values of $0.143$ or $0.5$ used in the cryo-EM literature [rosenthal_optimal_2003] are not meaningful in our plots showing FSC curves.
@@
(b) 1D cross-section of $\text{diag}(H)$
(c) $z=0$ slice of $H$
(d) Convergence of L-BFGS
Figure 3: (a) The condition number $\kappa(H)$ for the experimental dataset used in Section 4, with $N=30000$ images, grid of dimensions $192\times 192\times 192$, and $\lambda=10^{-8}$: lower bound of $\kappa(H)$ (blue, solid) computed using (21), the value of $\kappa(H)$ for the given dataset without CTF (red, dashed) and with CTF (green, dashdot). The condition numbers computed for this plot are for $H$ with nearest-neighbor interpolation. (b) The cross-section of the diagonal of the Hessian matrix $H$ computed in (a) without CTF (red, dashed) and with CTF (green, dashdot). (c) The $z=0$ plane of the diagonal of the Hessian matrix $H$ for the given dataset, computed using trilinear interpolation and reshaped as a 3D volume. Note the effect of the CTF, in contrast to Figure 1(c), where the particle images are not corrupted by the CTF. (d) The convergence of L-BFGS on problem (12) with trilinear interpolation and the dataset described in Section 4 to obtain the ground truth volume against which we evaluate the solutions of the SGD algorithms.
@@
Figure 4: Results of the numerical experiments with trilinear interpolation projection matrices $P_{\phi}^{tri}$ for $N=30000$ images of $192\times 192$ pixels. Panel (a) shows the loss function value and panel (b) shows the FSC of the final reconstruction with a ground truth solution for SGD with no preconditioner (blue), SGD with a precomputed preconditioner (red), and Algorithm 1, namely SGD with an estimated preconditioner (green). The black dashed line in (a) shows the loss function value of the final L-BFGS solution used as ground truth, against which we compute the FSC in (b).
@@
Ground truth (L-BFGS) No preconditioner Precomputed preconditioner Algorithm 1
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: A qualitative comparison of the reconstructions from the three SGD algorithms evaluated in Section 4 (columns 2-4) and the ground truth solution obtained using L-BFGS (first column), where each row shows the same view for all four volumes. Each volume is displayed as the isosurface at six standard deviations above its mean voxel value.
@@ 5 Outlook and conclusion
    A critical component of the current EM approaches is the $\ell_{2}$ regularizer, which makes the maximization step computationally tractable. SGD, on the other hand, is compatible with other regularization methods, and one could take advantage, for example, of the Wilson prior [gilles_molecular_2022] and learned regularization methods [kimanius_exploiting_2021, kamilov_plug-and-play_2023, kimanius_data-driven_2024].
@@ 6 Acknowledgments
The numerical experiments in their final form have been performed using the computational infrastructure at the MRC LMB; therefore, BT thanks Jake Grimmett, Toby Darling and Ivan Clayson of LMB Scientific Computing for computing support.
This work was supported by the NIH under grant R01GM136780, the Air Force Office of Scientific Research under grant FA9550-21-1-0317, the Alfred P. Sloan Foundation under grant FG-2023-20853, DARPA/DOD under grant HR00112490485, and the Simons Foundation. This research was performed while BT was affiliated with Yale University, with part of it while visiting the Institute for Pure and Applied Mathematics (IPAM), which is supported by the National Science Foundation (Grant No. DMS-1925919).
7 Data availability
The code for reproducing the numerical experiments is available on GitHub at
https://github.com/bogdantoader/simplecryoem and the particle metadata used for the numerical experiments as well as the output volumes and figures are available on Zenodo at
https://doi.org/10.5281/zenodo.14017756.